### v3

# Majority voting is not good for heaven or hell, with mirrored performance

Abstract
Within the ViSE (Voting in Stochastic Environment) model, we study the effectiveness of majority voting in various environments. By the pit of losses paradox, majority decisions in apparently hostile environments systematically reduce the capital of society. In such cases, the simple social decision rule of “rejecting all proposals without voting” outperforms majority. In this paper, we identify another pit of losses appearing in favorable environments. Here, the simple social decision rule of “accepting all proposals without voting” is superior to majority. We prove that under a version of simple majority called symmetrized majority and the antisymmetry of the voting body, the second pit of losses is a mirror image of the pit of losses in hostile environments and explain this phenomenon. Technically, we consider a voting society consisting of individualists whose strategy is supporting all proposals that increase their capital and a group (groups) whose members vote to increase the total group capital. According to the main result, the expected capital gain of each agent in the environment whose generator ${X}$ has mean $\mu>0$ exceeds by $\mu$ their expected capital gain under generator $-{X}$. This result extends to location families of generators with distributions symmetric about their mean. The mentioned result determines the symmetry of the difference between the expected capital gain under the symmetrized majority and that under the “basic” social decision rule that rejects/accepts all proposals in unfavorable/favorable environments.
keywords:
Voting paradoxes, Pit of losses, ViSE model, Majority voting, Symmetrized majority, Stochastic environment
pacs:
MSC Classification]91B70, 91B12, 91B14, 91B15, 90C40
1 Introduction
1.1 The problem of harmful voting
A large body of research on voting has led to the conclusion that this method of decision-making requires great care and attention when used. In particular, there are many voting paradoxes (see e.g., [36]), i.e., unexpected properties of voting, which in most cases are disappointing. A large class of paradoxes rests on a specific, carefully constructed agenda. The ViSE (Voting in Stochastic Environment) model introduces a stochastic agenda, however, this does not eliminate the paradoxical properties of majority decisions. Namely, the pit of losses paradox ([9, 26]) appears within the framework of this model. The essence of this paradox is that majority decisions in a markedly unfavorable environment are, on average, disadvantageous to society. In such cases, it is better to maintain the status quo by rejecting all proposals without a vote.
In this paper, we examine majority voting in favorable environments. It cannot, on average, reduce welfare. However, we demonstrate that just as outright rejection can be better than majority voting in unfavorable environments, total acceptance can outperform simple majority in a favorable environment. Moreover, under some conditions, these effects develop symmetrically along the favorability scale.
1.2 The ViSE model
The main assumptions of the ViSE model (for more details, see e.g., [9]) are as follows. A society consists of $n$ agents (also called voters). Each agent is characterized by the current real-valued capital (debt, if it is negative), which can sometimes be interpreted as utility. In each step $t=1,\ldots,T,$ some proposal is put to the vote, and the agents vote, guided by their voting strategies, for or against it. In the framework of the ViSE model, a strategy is an algorithm following which an agent uses available information to decide whether to vote for or against the proposal under consideration. In general, such a strategy may involve stochastic elements. Each proposal is a vector of algebraic capital gains of all agents. Proposals approved through the established social decision rule are implemented: the agents’ capitals change according to the proposal. If the proposal is rejected, the capitals do not change.
Each proposal is generated stochastically: it is a realization ${\mathbf{{x}}}=({{x}_{1}},\ldots,{{x}_{n}})={{\mathbf{{X}}}(\omega)},\,\omega\in\Omega$ of a multivariate random variable ${\mathbf{{X}}}=({{X}_{1}},\ldots,{{X}_{n}})$ called a proposal generator. In this notation, $\omega$ is an elementary outcome, $\Omega$ is a sample space. In expressions like ${\mathbb{E}}({{\mathbf{{X}}}(\omega)}),$ where ${\mathbb{E}}$ is the mathematical expectation operator, $\omega$ is assumed to range over $\Omega,$ so that ${\mathbb{E}}({{\mathbf{{X}}}(\omega)})\equiv{\mathbb{E}}({\mathbf{{X}}}).$ We consider the case where the components of ${\mathbf{{X}}}$ are independent and identically distributed with a given mean $\mu$ and standard deviation $\sigma$. The scalar random variable ${X}$ corresponding to the components of ${\mathbf{{X}}}$ will be called the gain generator.
The proposals put to the vote are called stochastic environment proposals. The environment is favorable if $\mu>0,$ neutral if $\mu=0,$ and unfavorable (hostile) if $\mu<0.$
The dynamics of agents’ capital in various environments can be analyzed to compare voting strategies and social decision rules in order to select the optimal ones in terms of maximizing appropriate criteria. Since in this version of the model the process is stationary, it can be described by one-step characteristics. An important one is the expected capital gain, ECG, of an agent with a certain strategy after implementing one collective decision of a given society in a given environment. The ECG (not to be confused with electrocardiogram) can be negative, zero, or positive.
In this paper, the environment is identified with the proposal generator ${\mathbf{{X}}}.$ Its properties influence the relationship between the current and future states of society. Thus, the model is applicable to situations where the issue is comparing the status quo with reform, rather than choosing among several candidates.
For an extended discussion of the ViSE model, its relation to reality, and research carried out within its framework, we refer to [25].
1.3 Relation between the ViSE model and other models
Dynamic models of voting in a multidimensional space of proposals have been intensively studied since the 1960s, see [32, 30, 37, 21]. However, proposal generation was usually the prerogative of the participants, either of the voters themselves (endogenous agenda), viz.,
  • (a)
    with the choice of a random proposer at every step ([3, 31, 20, 23]) or
  • (b)
    not completely randomly ([15, 11]),
or of some other agents with their own interests ([32, 34, 35, 39]).
Proposals generated exogenously and having random effects were considered in ([38, 12, 13]) [^1] In [13], the importance of analyzing dynamic voting in a stochastic environment was emphasized and the scarcity of relevant work was noted. The authors mentioned publications since 2008; the ViSE model ([7], etc.) differs significantly from the Dziuda and Loeper model, but belongs to the same class., but the non-transferable nature of the individual utilities (i.e., measuring utilities in subjective scales rather than in a common “currency”) made it difficult to study cooperative and prosocial strategies within the corresponding models. A simulation study of majority voting by bots with a random agenda in a model with ideal points was conducted in [6].
The studies of cooperation in dynamic voting in the game of pie sharing with transferable utilities started with [15]. An experimental study of cooperative solutions was performed in [14]. Interesting results on coalition dynamics and decision effectiveness in a model with discounting and bribes were obtained in [20]. The results of [9], where we looked for the optimal threshold of qualified majority, are comparable with the results of [24], where a stochastic model was studied for which, under certain conditions, simple majority was optimal. The problem of the optimal voting threshold was also considered in [10], where proposals were generated exogenously and the profitability of the first accepted proposal was evaluated; in that model, the unanimity rule would be optimal if there were no discounting.
A significant difference between the ViSE model and legislative bargaining models (such as a number of the models mentioned above, see also [4, 42]) is that looking for equilibria is the main problem of bargaining models, while the structure of the proposal space in the ViSE model, with the individual capitals as the coordinates, does not allow any equilibria in standard problem settings. This is caused by the lack of Pareto optimal alternatives in the state space, as the basic ViSE model allows simultaneous benefit of all agents in any state. In such circumstances, the goal becomes the effectiveness of decisions made, i.e., utility maximization: for certain categories of agents or for the whole society. Thereby the ViSE model provides a means to study the effectiveness of agents’ selfishness, group cooperation, and various prosocial strategies, as well as the effectiveness of social decision rules.
Thus, the ViSE model shifts the focus of research from finding equilibrium in bargaining to maximizing some aggregated utility, i.e., to the effectiveness under certain conditions.
The analysis of effectiveness brings this approach closer to the one implemented in [22], where there is no dynamics, since a number of (exogenously generated) proposals are voted “in parallel”, but the acceptance of any proposal yields some change in individual utilities, and the characteristics of the resulting gain vector allow us to evaluate the effectiveness of the decision-making mechanism.
In contrast to stochastic proposal models based on the Impartial Culture (IC) assumption and its variants, in the ViSE model the distribution of voters’ responses to a proposal depends on the properties of the environment.
It is worth noting that decision making within the ViSE model can be formalized as a general Markov decision process (see, e.g., [18, 41, 16]), where states are vectors of the current capital of the agents, and the possible social decision rules are as follows: vote by simple majority, vote using some other rules, accept the submitted proposal without voting, and reject the proposal without a vote. The last two rules can be treated as degenerate voting rules, namely, the threshold majorities with a negative threshold and with a threshold exceeding the number of voters, respectively. The probability (or density) of the transition between states $s$ and $s^{\prime}$ depends on the social decision rule and the probability (density) associated by the proposal generator with the proposal $s^{\prime}-s$. The society’s reward is the sum of the components of an accepted proposal, or zero if the proposal is rejected. This perspective is useful for interpreting research results within the ViSE model.
1.4 Pit of losses paradox
It is well known that agenda control allows, through a series of votes, to bring a voting society into a state worse than the initial state for all voters. Moreover, in most cases, the society can be brought to any state. This holds true for various models of sequential voting in a multidimensional policy space ([27, 28, 29, 32]) and is known as the so-called chaos theorems in social choice theory.
This is an important and paradoxical knowledge about social reality. It refutes the idea a la Adam Smith that a society composed of economic agents driven by self-interest will prosper through the invisible hand of democracy. No, through agenda control you can do anything you want to this society, even make it self-destruct. The majority’s gain may turn into the loss of all. There may be hope, however, that the issue is agenda manipulation, and by excluding it this effect can be eliminated.
The ViSE model leaves no room for a purposeful agenda: it is replaced by a favorable or unfavorable stochastic environment that generates proposals. However, a voting paradox of the above nature still arises. Namely, by [9, Proposition 1], if $\mu,$ the mean of a Gaussian gain generator ${X}$, is below a certain negative threshold (which is proportional to the standard deviation of ${X}$ and depends on the number of voters), then for a purely individualistic society and majority rule, the expected reward (one-step capital gain, ECG) of each agent is also negative. This is the pit of losses paradox illustrated in Fig. 1a. It means that sequential voting will bankrupt all participants in such a society.
(a)                                                      (b)
Figure 1: The expected one-step capital gain (ECG) of an agent in a voting society of 21 individualists: (a) in unfavorable environments $(\mu<0)$; (b) in some favorable environments $(\mu>0)$
This implies that in some unfavorable environments, majority rule is worse for each agent than rejecting all proposals without a vote, which saves capital. For gain generators distributed with longer tails than Gaussian generators, the pit of losses also exists, but it is shallower.
Thus, the essence of the paradox contained in the chaos theorems [^2] It should be noted, however, that sometimes decision making routines that encourage chaotic conflict are considered to be robust to bounded rationality and “promote effective search in an uncertain task environment” ([19]). Furthermore, chaos is not absolute as under a random agenda some areas are more attractive than others (see, e.g., [6]). is not agenda manipulation, but the combination of selfishness with majority rule. Methods to resolve the pit of losses paradox include cooperation ([8]), prosocial behavior ([43, 44]) that can be supported by income redistribution through taxes ([1]), and optimizing the voting threshold ([9, 26]; see Section 4 below).
1.5 Contributions of this study
The pit of losses paradox is characteristic of unfavorable environments. What happens in favorable environments? One might assume that the agent’s expected capital gain curve approaches the line $\mathrm{ECG}=\mu$ in the first quadrant from above as $\mu$ increases, i.e., that majority decisions provide some decreasing positive gain over what the environment offers.
However, somewhat surprisingly, this curve intersects the the line $\mathrm{ECG}=\mu$ (or $\mathrm{ECG}/\sigma=\mu/\sigma$, as shown in Fig. 1b). This means that under sufficiently favorable conditions, filtering proposals by majority vote is harmful. Furthermore, the abscissa of the intersection is equal in absolute value to the zero of the same function, marked with a black circle in Fig. 1a. In Sections 3 and 4 we study and explain this phenomenon in the more general case of a society consisting of individualists and a group (or groups).
Let us consider the following agent strategies. An individualist casts a vote (half a vote) for a proposal whenever this proposal increases (respectively, preserves) their own capital. The ‘half-vote’ outcome is provided to ensure the complementarity effect of opposite proposals in the case of discrete/discontinuous distributions. The voting strategy of a group member is to cast a vote (half a vote) for a proposal whenever this proposal increases (respectively, preserves) the total capital of the group.
The question under consideration is: “How effective is majority voting compared to the simplest decision-making rules (such as unconditional acceptance or rejection) in the environments of various favorability?”
Lemma 1 below shows that under some conditions and any generator ${\mathbf{{X}}}$ with mean $\mu>0$, the expected capital gain of any agent exceeds that under generator $-{\mathbf{{X}}}$ by $\mu.$ In Theorem 3 and Remark 1, this result is applied to voting bodies consisting of individualists and group members. Theorem 4, an easy consequence of the previous one, states that the same difference of $\mu$ holds for generators ${\mathbf{{X}}}^{\mu}$ with mean $\mu$ and ${\mathbf{{X}}}^{-\mu}={\mathbf{{X}}}^{\mu}-2\mu,$ provided that ${\mathbf{{X}}}^{\mu}$ is symmetric about $\mu.$ In Section 4 we introduce the concept of expected relative capital gain (ERCG) and show that under the conditions of Theorem 4, the ERCG of any agent is an even function of $\mu,$ i.e., $\mbox{ERCG}(-\mu)=\mbox{ERCG}(\mu)$ for any $\mu$. This allows us to conclude that the pit of loses paradox extends to favorable environments, and the performance of the symmetric majority rule is symmetric about 00. This is why “majority voting is not good for heaven or hell, with mirrored performance” (see Section 4).
Thus, this study contributes to the literature on the peculiarities and limitations of majority voting (see, e.g., [45, 5, 36, 2, 40, 33]). More specifically, it identifies some of such limitations in the context of varying favorability of the environment, which is the hallmark of research within the ViSE model.
2 Notation and problem formulation
We consider a society in which $\ell$ agents are individualists, while $g$ agents form a group, so that $g+\ell=n$ is the size of the society.
Let us say that any $n$ agents with their voting strategies and an adopted social decision rule form a voting body.
The social decision rule we consider is simple majority (a proposal is rejected/accepted if it receives no more/more than $n/2$ votes) modified by means of the random tie-breaking mechanism that accepts a proposal with probability $1/2$ when exactly $n/2$ votes are cast for it (cf., e.g., [17]). This rule will be called the symmetrized majority rule. When the gain generator is continuous and the number of agents is odd, this rule is equivalent to the simple majority rule, and we will use the latter name.
We say that a voting body is antisymmetric if for any proposal ${\mathbf{{x}}},$ the probabilities of ${\mathbf{{x}}}$ and $-{\mathbf{{x}}}$ being accepted by that voting body sum to $1$. An interpretation of this is that a voting body makes complementary stochastic decisions regarding any opposite proposals.
Given a voting body, let $n^{+}({\mathbf{{x}}})$ be the number of votes cast for proposal ${\mathbf{{x}}}$; let $I({\mathbf{{x}}})$ be the probability of accepting (and thus implementing) ${\mathbf{{x}}}$. If the social decision rule is not stochastic, then for any proposal ${\mathbf{{x}}},$ $I({\mathbf{{x}}})$ is either 00 or $1$
Let ${\mathbf{D}}_{\mathbf{{X}}}$ denote the random vector of capital gains that implements the collective decision on the proposal generated by ${\mathbf{{X}}}.$
In short, $I({\mathbf{{x}}})$ is the implementation probability of proposal ${\mathbf{{x}}}$; ${\mathbf{D}}_{\mathbf{{X}}}$ is the vector of implemented capital gains, where ${\mathbf{{X}}}$ is the random vector of proposed capital gains.
Thus, the main notations (articles omitted) are as follows:
$\ell$ number of individualists
$g$ size of the group
$n=g+\ell$ number of agents
${\mathbb{E}}(\mathbf{Z})$ multivariate mean of random vector $\mathbf{Z}$
${X}={X}^{\mu}$ gain generator having a mean $\mu$ and a finite variance
$\mu=\mu\mathop{\!{}_{{X}}}={\mathbb{E}}({X})$ mean of ${X}$
$\sigma=\sigma\mathop{\!{}_{{X}}}$ standard deviation of ${X}$
${\mathbf{{X}}}={\mathbf{{X}}}^{\mu}$ proposal generator corresponding to ${X}^{\mu}$
${\mathbf{D}}_{\mathbf{{X}}}$ row vector of implemented capital gains for generator ${\mathbf{{X}}}$
${\mathbf{{X}}}=({{X}_{1}},\ldots,{{X}_{n}})$ generator ${\mathbf{{X}}}$ with ${{X}_{1}},\ldots,{{X}_{g}}$ corresponding to the group
${\mathbf{{x}}}=({{x}_{1}},\ldots,{{x}_{n}})$ proposal, a realization of proposal generator ${\mathbf{{X}}}$
$n^{+}({\mathbf{{x}}})$ number of votes cast for proposal ${\mathbf{{x}}}$
$I({\mathbf{{x}}})$ implementation probability of proposal ${\mathbf{{x}}}$
The purpose of this study is to find out how the effectiveness of majority decisions depends on the favorability of the environment. In particular, we examine the relationship between ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}_{1}}\bigr{)}$ and ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}_{2}}\bigr{)}$ for arbitrary opposite proposal generators ${\mathbf{{X}}}_{1}$ and ${\mathbf{{X}}}_{2}$, with two different notions of opposition.
3 Main results
3.1 Relationship between gains under opposite proposals
The following lemma establishes a connection between the implemented capital gains under proposal generators ${\mathbf{{X}}}={\mathbf{{X}}}^{\mu}$ and $-{\mathbf{{X}}}$ (such generators will be called opposite ones) when a voting body is antisymmetric.
Lemma 1.
For any gain generator ${X}$ with mean $\mu$ and any antisymmetric voting body, it holds that
${\mathbb{E}}({\mathbf{D}}_{-{\mathbf{{X}}}})={\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}})-\mu\bm{1},$
where $\bm{1}=(\underbrace{1,\ldots,1}_{n})$.
Proof of Lemma 1.
Let ${\mathbf{{Y\!}}_{\mathbf{{X}}}}$ be the function of proposal generator ${\mathbf{{X}}}$ defined by
${\mathbf{{Y\!}}_{\mathbf{{X}}}}(\omega)={{\mathbf{{X}}}(\omega)}\hskip 1.1ptI({{\mathbf{{X}}}(\omega)}),\quad\omega\in\Omega.$ (1)
By the definition of ${\mathbf{D}}_{\mathbf{{X}}},$ for any $\omega\in\Omega,$ we have
${\mathbb{E}}\bigl{(}{\mathbf{D}}_{\mathbf{{X}}}(\omega)\,|\>\omega\bigr{)}={{\mathbf{{X}}}(\omega)}\hskip 1.1ptI({{\mathbf{{X}}}(\omega)})+0\cdot\!(1-I({{\mathbf{{X}}}(\omega)}))={\mathbf{{Y\!}}_{\mathbf{{X}}}}(\omega).$
The law of total expectation yields
${\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}})={\mathbb{E}}({\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}}(\omega)\,|\>\omega))={\mathbb{E}}({\mathbf{{Y\!}}_{\mathbf{{X}}}}).$ (2)
Since the voting body is antisymmetric,
[…]
By (1), (2), and (3), for any ${\mathbf{{X}}}$, it holds that
[…]
$=$ ${\mathbb{E}}\bigl{(}-{{\mathbf{{X}}}(\omega)}(1-I({{\mathbf{{X}}}(\omega)}))\bigr{)}={\mathbb{E}}(-{{\mathbf{{X}}}(\omega)})+{\mathbb{E}}({{\mathbf{{X}}}(\omega)}\hskip 1.1ptI({{\mathbf{{X}}}(\omega)}))$
$=$ $-{\mathbb{E}}({\mathbf{{X}}})+{\mathbb{E}}\bigl{(}{\mathbf{{Y\!}}_{\mathbf{{X}}}}(\omega)\bigr{)}=-\mu\bm{1}+{\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}}).$
This completes the proof. ∎
Lemma 2.
Any voting body formed by individualists, members of one group, and the symmetrized majority rule is antisymmetric.
Proof of Lemma 2.
Observe that $\tfrac{1}{2}(\mathop{\rm sgn}\nolimits{{x}_{i}}+1)$ is the number of votes cast by individualist $i$ for proposal ${\mathbf{{x}}}=({{x}_{1}},\ldots,{{x}_{n}}).$ Similarly, $\tfrac{1}{2}\left(\mathop{\rm sgn}\nolimits\sum_{i=1}^{g}{{x}_{i}}+1\right)$ is the number of votes cast by each group member. Then for any voting body of the given type and any ${\mathbf{{x}}}={{\mathbf{{X}}}(\omega)},\,\omega\in\Omega$, we have
$n^{+}({\mathbf{{x}}})+n^{+}(-{\mathbf{{x}}})$ $=$ $\tfrac{1}{2}\left(\mathop{\rm sgn}\nolimits\sum_{i=1}^{g}{{x}_{i}}+1\right)g+\tfrac{1}{2}\!\sum_{i=g+1}^{n}(\mathop{\rm sgn}\nolimits{{x}_{i}}+1)$
$+$ $\tfrac{1}{2}\left(\mathop{\rm sgn}\nolimits\sum_{i=1}^{g}(-{{x}_{i}})+1\right)g+\tfrac{1}{2}\!\sum_{i=g+1}^{n}(\mathop{\rm sgn}\nolimits(-{{x}_{i}})+1)$
$=$ $g+\ell=n.$
Therefore, matching the corresponding relations for $n^{+}({\mathbf{{x}}})$ and $n^{+}(-{\mathbf{{x}}})$ we have
$n^{+}({\mathbf{{x}}})>(=)\,(<)\;\tfrac{n}{2}\,\Leftrightarrow\,n^{+}(-{\mathbf{{x}}})<(=)\,(>)\;\tfrac{n}{2}\,.$ (4)
Since by the definition of symmetrized majority,
$n^{+}({\mathbf{{x}}})>(=)\,(<)\;\tfrac{n}{2}\,\Leftrightarrow\,I({\mathbf{{x}}})=1\,\bigl{(}\tfrac{1}{2}\bigr{)}\,(0),$
it follows from (4) that $I({\mathbf{{x}}})+I(-{\mathbf{{x}}})=1$. Thus, the voting body is antisymmetric. ∎
Theorem 3.
For any gain generator ${X}$ with mean $\mu$ and for any voting body formed by individualists, members of one group, and the symmetrized majority rule, it holds that
${\mathbb{E}}({\mathbf{D}}_{-{\mathbf{{X}}}})={\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}})-\mu\bm{1}.$
Theorem 3 follows from Lemmas 1 and 2.
Remark 1.
Similarly, it can be shown that Theorem 3 extends to all voting bodies consisting of individualists, any number of groups, and the symmetrized majority rule. Moreover, each group may vote either to increase the total group capital ((((as previously assumed)))) or under the intra-group symmetrized majority rule. In the latter case, every group member supports the decision made by the intra-group symmetrized majority, and the group is called an alternative one (while the term “group” is retained for groups of the former type). Since an individualist is a special case of a group of either type, the formulation reduces to: Theorem 3 extends to all voting bodies formed by any number of groups of arbitrary types and the symmetrized majority rule.
3.2 Gains on location families of generators
For any ${X}^{0},$ which is a gain generator with mean $0,$ consider the location family of gain generators
${X}^{\mu}={X}^{0}+\,\mu,\quad\mu\in{\mathbb{R}}.$ (5)
For such families, the following theorem holds, which requires the symmetry of the gain distribution around its mean.
Theorem 4.
Suppose that society consists of individualists and a group, the social decision rule is symmetrized majority, and the distribution of the gain generator ${X}^{\mu}$ is symmetric about $\mu.$ Then
${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}^{-\!\mu}}\bigr{)}={\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}^{\mu}}\bigr{)}-\mu\bm{1},$ (6)
where ${\mathbf{D}}_{{\mathbf{{X}}}^{\mu}}$ is the vector of implemented capital gains under ${X}^{\mu}$.
Proof.
Since the distribution of ${X}^{\mu}$ is symmetric about $\mu$, gain generators $-{X}^{\mu}$ and ${X}^{-\mu}={X}^{\mu}-2\mu$ share the same distribution. Therefore, the desired statement follows from Theorem 3. ∎
In the general case of not necessarily symmetric distributions, generators ${\mathbf{{X}}}^{\mu}$ and ${\mathbf{{X}}}^{-\mu}$ can be referred to as quasi-opposite ones and (6) is not guaranteed.
Like Theorem 3, Theorem 4 can be extended as described in Remark 1.
4 Interpretation and discussion
Consider the following social decision rule:
Reject all proposals if $\mu\leq 0;$ accept all proposals if $\mu>0$.
This rule will be referred to as the basic rule.
For any social decision rule $\psi,$ let the expected relative capital gain (ERCG) of an agent under $\psi$ be the mathematical expectation of the difference between the agent’s capital gain under $\psi$ and that under the basic rule.
For families (5), consider the ERCG as a function of $\mu.$
Corollary 5.
[…]
Proof.
For any family (5) and any agent, the expected capital gain under the basic rule is 00 when $\mu\leq 0$ and $\mu$ when $\mu>0.$ Therefore, for any $\mu>0,$ the left-hand and right-hand sides of (6) are the vectors of agents’ expected relative capital gains under gain generators ${X}^{-\mu}={X}^{\mu}-2\mu$ and ${X}^{\mu}$, respectively. Since by Theorem 4 they are equal, the ERCG is an even function of $\mu.$
(a)..(a)                                                      (b)
Figure 2: Expected relative capital gain: (a) in the society of 21 individualists under majority voting (red) and voting with the optimal threshold (green); (b) in the society consisting of a group of 7 agents and 14 individualists (for comparison, the ERCG curves for homogeneous societies are added)
For the society of 21 individualists with Gaussian generators (5), the expected capital gain under simple majority rule was shown in Fig. 1. The ERCG for this family is presented in Fig. 2a. In addition to a pit of losses in unfavorable environments, this function has a mirror image of that pit in favorable environments. Decisions made by majority votes in these two areas are, on average, disadvantageous to society compared to the basic rule.
The reason for the pit of losses paradox is that in unfavorable environments, the average positive gain is smaller, while the average absolute loss is larger than in the neutral environment of the same variance (where these average values are equal). Therefore, the average difference between [the absolute values of] a gain and a loss is negative. Consequently, the proposals accepted by simple majority with a small margin of votes in favor (which is typical in an unfavorable environment) yield, on average, a total loss. Thus, in such environments, the majority’s gain does not compensate the loss suffered by the minority. Note that this imbalance identified using the ViSE model is relevant to practice.
The nature of the positive-area pit of losses is dual: while majorities typically approve harmful (with respect to the total capital) proposals in unfavorable environments, under favorable conditions they quite often reject proposals that provide society with positive total gains.
Fig. 2a allows us to conclude that if the set of environments with $\mu/\sigma<-0.254$ is considered as “hell” and the set of environments with $\mu/\sigma>0.254$ as “heaven”, then the simple majority rule is worse than the basic rule in both of them. Simple majority outperforms the basic rule only in a vicinity of the neutral environment ($\mu=0$), where the ERCG peak is quite high.
For the society with a group of 7 agents and 14 individualists (Fig. 2b), Gaussian gain generators produce a completely positive ERCG for group members (blue curve). In contrast, the ERCG of individualists is negative, except for a narrow interval of $\mu/\sigma\in[-0.03,0.03]$ (red curve).
It is instructive to compare this with the ERCG in the societies of 21 individualists (red dotted curve) and of 21 group members (blue dotted curve). The most advantageous position is in the group of 7 agents and the worst is among the individualists complementing this group. Such a group can be considered as an “elite”. It usually receives relatively high capital gains at the expense of individualists. It should be noted that the ERCG of the whole mixed society (gray curve) is even lower than that for the society of 21 individualists. The impact of a “responsible elite” on public welfare was investigated in [43, 44]. In most cases, such an elite provides society with capital growth.
Obviously, the most beneficial strategy for the whole society is to be one cohesive group. However, in practice this is unrealistic. While a group of like-minded people is able to develop a common objective function, this is problematic for society. Therefore, non-consensus-based decisions are of primary interest. The “basic rule”, which outperforms the simple majority in both “hell” and “heaven” is extremely crude. How can taking into account voters’ opinions improve the outcome? Such a social decision rule is voting with the optimal threshold ([9, 26]). This threshold is exactly $0.5$ in the neutral environment, as well as for $\mu/\sigma\in[-0.1,0.1]$ in our example (Fig. 2a). Outside a certain interval around 00, the optimal threshold is usually (but not always) higher in unfavorable and lower in favorable environments. The ERCG curve in the case of voting with the optimal threshold clearly has no pit of losses. For the society of 21 individualists, it is shown in Fig. 2a.
5 Conclusion
The ViSE model is designed to evaluate the effectiveness of individual voting strategies and social decision rules in environments of varying favorability and, more generally, in environments with various gain distributions. In this paper, we show that in the case of Gaussian environments and two-component (a group and individualists) societies, the symmetrized majority rule is effective only under more or less neutral conditions, while in “heaven” or “hell” it is harmful. Furthermore, for environments with symmetric distributions, the performance of this rule is symmetric about the neutrality point. For general distributions with finite mean, the performance is the same for any two opposite generators.
[…]
Acknowledgments We are very grateful to two anonymous referees, especially, for suggesting the idea of reverse time to obtain a simpler proof of the technical result. In the final version, we use a very similar approach in other terms, supplemented by the condition of antisymmetry of the voting body, which is necessary to generalize the result. We also thank the participants of the seminar on stochastic voting models at the Institute of Control Sciences of the Russian Academy of Sciences: Yana Tsodikova, Anton Loginov, Vitaly Malyshev, Vladislav Maksimov, Anna Khmelnitskaya, Irina Kalushina, Zoya Lezina, Vladimir Borzenko.
Compliance with ethical standards
Funding   Work of P.C. was supported by the European Union (ERC, GENERALIZATION, 101039692). Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Conflict of interest   The authors have no relevant financial or non-financial interests to disclose that are relevant to the content of this article and declare that they have no conflict of interest.
Ethical approval   This article does not contain any studies with human participants or animals performed by any of the authors.
Consent to participate   Not applicable.
Consent for publication   The authors gave explicit consent to publish this manuscript.
Availability of data and materials   Not applicable.
Code availability   Not applicable.